OpenStudy (anonymous):

a garden has the shape of a right triangle with one leg 3 meters longer than the other. the hypotenuse is 3 meters less than twice the length of the shorter leg. what is the length of the shorter leg? Please help!

OpenStudy (anonymous):

If the shortest leg is x, the side at right angles to it is x+3 and the hypotenuse is 2x-3 Using Pythagoras x^2 + (x+3)^2 = (2x-3)^2 x^2 + x^2 +6x + 9 = 4x^2 - 12x +9 2x^2 + 6x + 9 = 4x^2 - 12x +9 2x^2 -18x = 0 2x(x - 9) = 0 either 2x = 0 or x-9 = 0 x = 0 or x = 9